# A Gentle Wakeup Call Symmetry Breaking with Less Collision Cost

### v2

@@ 4 Upper Bound
The next lemma establishes bounds on expected latency and total collision cost, and the intuition is as follows. In Case 2, collisions are “cheap” by Lemma 5, in the sense that $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, the cost of collisions incurred until the good window is reached is small; it takes only $\log_{2}(n\sqrt{\mathcal{C}})+O(1)=O(\log n)$ doublings of the window size to reach a good window in the doubling phase. Since $2^{j}$ counts the number of doublings we perform in the doubling phase (recall Section 3), we reach a good window by $j=O(\log\log n)$. After this point the probability of success is high, and so subsequent iterations contribute little to the expected total cost of collisions.
Lemma 6
@@
Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, the expected latency is $O\left(\log^{1/(2\epsilon)+4}n)\right)$.
Proof
@@
The number of slots incurred for each individual halving phase is the sample size times the number of halvings, so upper bounded by:
$O\left(\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\right)\lg(w_{0})$ $=O\left(\sqrt{\mathcal{C}}\log(2^{j}w_{0})\log(w_{0})\right)$
$=O\left(\sqrt{\mathcal{C}}j\mathcal{C}^{2\epsilon}\right)$
$=O\left(j\log^{1/(2\epsilon)+2}n\right)$
@@
where the last line follows by applying Lemma 5. Thus, over iterations $1$ to $j$, the number of slots contributed by the halving phases is $O(j^{2}\log^{1/(2\epsilon)+2}n)$
The number of slots in the doubling phase of the $j$-th iteration is:
$2^{j}d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ $=O(2^{j}\sqrt{\mathcal{C}}\log(2^{j}w_{0}))$
$=O(2^{j}\sqrt{\mathcal{C}}\,j\mathcal{C}^{\epsilon})$
$=O(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n))$
@@
slots, where the last line follow by Lemma 5. By the above, summing up the number of slots executed over iterations 1 to $j$ is $O(2^{j}\,j^{2}\,(\log^{1/(2\epsilon)+2}n))$.
@@
Pessimistically, let us assume no success ever occurs during the halving phase. Focusing on the doubling phase, a good window is reached within at most $\log_{2}(n\sqrt{\mathcal{C}})+O(1)=\log_{2}n+O(\log\log n)$ doublings of the window size. Since $2^{j}$ counts the number of doublings we perform in the doubling phase (recall Section 3), we reach a good window by iteration $j=c^{\prime}\log\log n$, where $c^{\prime}>0$ is a sufficiently large constant. By the above, the total number of slots up until a good window is reached is:
$O\left(2^{j}\,j^{2}\,(\log^{1/(2\epsilon)+2}n)\right)$ $=O\left((\log n)\,(\log\log n)^{2}(\log^{1/(2\epsilon)+2}n)\right)$
$=O\left(\log^{1/(2\epsilon)+4}n\right)$
@@
After this point the probability of success is high, by Lemma 1, and so subsequent iterations contribute little to the expected total cost of collisions. The expected latency is:
$O\left(\log^{1/(2\epsilon)+4}n\right)+\hskip-8.0pt\sum_{j=c^{\prime}\lg\lg n}^{\infty}\hskip-8.0ptO\left(\frac{1}{\eta}\right)^{j-c^{\prime}\lg\lg(n)+1}\hskip-5.0ptO\left(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n)\right)$
$=O\left(\log^{1/(2\epsilon)+4}n\right)+\hskip-8.0pt\sum_{j=c^{\prime}\lg\lg n}^{\infty}\hskip-8.0ptO\left(\frac{1}{n}\right)^{j-c^{\prime}\lg\lg(n)+1}\hskip-10.0ptO\left(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n)\right)$
$=O\left(\log^{1/(2\epsilon)+4}n\right)$
which completes the argument.
@@
The following lemma about the expected total cost of collisions is proved similarly to Lemma 6. The main difference is that, even if every slot contains a collision, the cost increases only by a $\mathcal{C}=O(\log^{1/\epsilon}n)$ factor.
Lemma 7
@@
Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, the expected total cost of collisions is $O\left(\log^{3/(2\epsilon)+4}n\right)$.
4.3 Performance of AH
We can now prove our upper bounds on the latency and expected collision cost for AH.
@@ 4.3 Performance of AH
  2. 2.
    Else, the expected latency is $O(\log^{4+\frac{1}{2\epsilon}}n)$ and the expected total cost of collisions is $O(\log^{4+\frac{3}{2\epsilon})}n)$.
Proof
The relationship between $n$ and $\mathcal{C}$ is established by the adversary (recall Section 1.1), and there are there are two cases to consider. In Case 1, $n\sqrt{\mathcal{C}}\leq w_{0}$. By Lemmas 3 and 4, the expected latency and the expected total collision cost for AH are each $O(\mathcal{C}^{(1/2)+2\epsilon})$, yielding (1).
In the second case, $n\sqrt{\mathcal{C}}>w_{0}$. By Lemmas, 6 and 7, the expected latency and expected total cost of collisions are $O(\log^{4+\frac{1}{2\epsilon}}n)$ and $O(\log^{4+\frac{3}{2\epsilon})}n)$, respectively, yielding (2).
5 Lower Bound
@@
$TC(\mathcal{A})=\sum_{u}LC(u)=n\cdot LC(u)$
By our assumption, $n\cdot LC(u)<2$. The probability of $u$ not succeeding in slot any fixed $t$ is at least 1p(t)1-p(t). Therefore, the probability that $u$ never succeeds over the execution of $\mathcal{A}$ is:
$\geq\prod_{i=1}^{\lambda}(1-p(i))$
$\geq e^{-2\sum_{i}^{\lambda}p(i)}$
$\geq e^{-2\cdot LC(u)}$
$\geq e^{-2\cdot TC(\mathcal{A})/n}$
$>e^{-4/n}$
where the second line follows from Fact 1(b), and the third line follows from the definition of $LC(u)$. The last line follows from the assumption that $TC(\mathcal{A})<2$.
It follows that the probability of all packets failing over the execution of $\mathcal{A}$ is:
$\geq\left(e^{-4/n}\right)^{n}=\frac{1}{e^{4}}$
Thus, there is a constant probability of all packets failing, which contradicts the w.h.p. guarantee of $\mathcal{A}$.
[…]
Lemma 9(Biswas et al. [5], Lemma 24)
@@
By Lemma 9, any algorithm that has a slot with contention higher than $2$ has an expected collision cost that is $\Omega(\mathcal{C})$. Therefore, we need only to consider algorithms where every slot in the execution has contention at most $2$. Under these circumstances, we provide a lemma that lower bounds the probability of a collision as a function of contention. We use $\texttt{Con}^{2}(t)$ to denote $(\texttt{Con}(t))^{2}$.
Lemma 10
@@
Proof
For a fair algorithm, in any slot $t$, $\texttt{Con}(t)=np$. For any fixed slot, the probability of a collision involving exactly two packets (since there are at least $2$ packets by assumption) is a lower bound on the probability of a collision (that may involve more than $2$ packets), and this is $\binom{n}{2}p^{2}(1-p)^{n-2}$, which is at least:
$\geq\binom{n}{2}p^{2}(1-p)^{n-2}$
$\geq\binom{n}{2}p^{2}(1-p)^{n}\geq\left(\frac{n^{2}}{2}\right)\left(\frac{\texttt{Con}(t)}{n}\right)^{2}e^{-2\cdot\texttt{Con}(t)}\geq\left(\frac{\texttt{Con}^{2}(t)}{2e^{2\cdot\texttt{Con}(t)}}\right)$
$=\Omega(\texttt{Con}^{2}(t))$
where the last line follows from $\texttt{Con}(t)\leq 2$.
@@
Proof
Let $\mathcal{A}$ be an algorithm that w.h.p. in $\eta$ solves wakeup with $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$ latency, where $d^{\prime}>0$ is any constant. By Lemma 8, $\mathcal{A}$ has total contention:
$\sum_{t=1}^{L}\texttt{Con}(t)\geq 2$ (6)
As we discussed, an implication of Lemma 9, we need only consider that $\mathcal{A}$ has contention at most $2$ in every slot. Thus, we can use Lemma 10 (noting that $n\geq 2$), which implies that the expected cost of collisions in a single slot $t$ is:
@@
$\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)\mathcal{C}$
By Jensen’s inequality for convex functions, we have:
$\frac{\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)}{L}\geq\left(\frac{\sum_{t=1}^{L}k\,\texttt{Con}(t)}{L}\right)^{2}$
which implies:
$\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)\geq\frac{\left(\sum_{t=1}^{L}k\,\texttt{Con}(t)\right)^{2}}{L}$ (7)
So, by Equation 7, the expected cost is at least:
[…]
and since $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$, we have:
$\frac{\mathcal{C}\,k^{2}\,2^{2}}{L}$ $=\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$
which yields the claim.
@@
6 Conclusion and Future Work
The theoretical exploration of collision costs remains in its early stages. Our work here provides steps towards addressing fundamental problems that connect with the wireless domain.
For future work, we aim to prove more general lower bounds. Additionally, we wish to derive results for the dynamic setting, where packets arrive over time rather than all being initially present.

### v3

@@ 4. Upper Bound
The next lemma establishes bounds on expected latency and total collision cost, and the intuition is as follows. In Case 2, collisions are “cheap” by Lemma 4.5, in the sense that $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, the cost of collisions incurred until the good window is reached is small; it takes only $O(\log n)$ doublings in the doubling phase. After this point the probability of success is high, and so subsequent iterations contribute little to the expected total cost of collisions.
Lemma 4.6.
@@
Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, the expected latency is $O\left(\log^{1/(2\epsilon)+5}n)\right)$.
Proof.
@@
The number of slots incurred for each individual halving phase is the sample size times the number of halvings, which is:
$O\left(\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\right)\lg(w_{0})$ $=O\left(\sqrt{\mathcal{C}}\log(w_{0})\log(w_{0})\right)$
$=O\left(\sqrt{\mathcal{C}}\mathcal{C}^{2\epsilon}\right)$
$=O\left(\log^{1/(2\epsilon)+2}n\right)$
@@
where the second line follows from $w_{0}=2^{\mathcal{C}^{\epsilon}}$, and the last line follows by applying Lemma 4.5. Thus, over iterations $1$ to $j$, the number of slots contributed by the halving phases is $O(j^{2}\log^{1/(2\epsilon)+2}n)$
The number of slots in the doubling phase of the $j$-th iteration is at most:
$jd\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ $=O\left(j\sqrt{\mathcal{C}}\log(2^{j}w_{0})\right)$
$=O\left(j\sqrt{\mathcal{C}}\,j\mathcal{C}^{\epsilon}\right)$
$=O\left(j^{2}\,(\log^{1/(2\epsilon)+1}n)\right)$
@@
slots, where the last line follow by Lemma 4.5. By the above, summing up the number of slots executed in the doubling phases over iterations 1 to $j$ is $O(j^{3}\,(\log^{1/(2\epsilon)+2}n))$.
@@
Pessimistically, let us assume no success ever occurs during the halving phase. Focusing on the doubling phase, a good window is reached within at most $\log_{2}(n\sqrt{\mathcal{C}})+O(1)=\log_{2}n+O(\log\log n)$ doublings of the window size. Since $j$ counts the number of doublings we perform in the doubling phase (recall Section 3), we reach a good window by iteration $j=c^{\prime}\log n$, where $c^{\prime}>0$ is a sufficiently large constant. By the above, the total number of slots up until a good window is reached is:
$O\left(j^{3}\,(\log^{1/(2\epsilon)+2}n)\right)$ $=O\left((\log^{3}n)(\log^{1/(2\epsilon)+2}n)\right)$
$=O\left(\log^{1/(2\epsilon)+5}n\right)$
@@
After this point the probability of success is high, by Lemma 4.1, and so subsequent iterations contribute little to the expected total cost of collisions. The expected latency is:
$O\left(\log^{1/(2\epsilon)+5}n\right)+\hskip-8.0pt\sum_{j=c^{\prime}\lg\lg n}^{\infty}\hskip-8.0ptO\left(\frac{1}{\eta}\right)^{j-c^{\prime}\lg\lg(n)+1}\hskip-5.0ptO\left(j^{3}\,(\log^{1/(2\epsilon)+2}n)\right)$
$=O\left(\log^{1/(2\epsilon)+5}n\right)$
which completes the argument. ∎
@@
The following lemma is proved similarly to Lemma 4.6. The main difference is that, even if every slot contains a collision, the cost increases only by a $\mathcal{C}=O(\log^{1/\epsilon}n)$ factor.
Lemma 4.7.
@@
Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, the expected total cost of collisions is $O\left(\log^{3/(2\epsilon)+5}n\right)$.
4.3. Performance of AH
We can now prove our upper bound.
@@ 4.3. Performance of AH
  2. (2)
    Else, the expected latency is $O(\log^{5+\frac{1}{2\epsilon}}n)$ and the expected total cost of collisions is $O(\log^{5+\frac{3}{2\epsilon}}n)$.
Proof.
The relationship between $n$ and $\mathcal{C}$ is established by the adversary (recall Section 1.1), and there are two cases to consider. In Case 1, $n\sqrt{\mathcal{C}}\leq w_{0}$. By Lemmas 4.3 and 4.4, the expected latency and the expected total collision cost for AH are respectively $O(\mathcal{C}^{(1/2)+2\epsilon})$ and $O(\mathcal{C}^{(1/2)+\epsilon})$, yielding (1).
In the second case, $n\sqrt{\mathcal{C}}>w_{0}$. By Lemmas, 4.6 and 4.7, the expected latency and expected total cost of collisions are $O(\log^{5+\frac{1}{2\epsilon}}n)$ and $O(\log^{5+\frac{3}{2\epsilon}}n)$, respectively, yielding (2). ∎
5. Lower Bound
@@
Lemma 5.2 (Biswas et al. (Biswas et al., 2024b), Lemma 24).
@@
By Lemma 5.2, any algorithm that has a slot with contention higher than $2$ has an expected collision cost that is $\Omega(\mathcal{C})$. Therefore, we need only to consider algorithms where every slot in the execution has contention at most $2$. Under these circumstances, we provide a lemma that lower bounds the probability of a collision as a function of contention. We use $\texttt{Con}^{2}(t)$ to denote $(\texttt{Con}(t))^{2}$.
Lemma 5.3.
@@
Proof.
Let $\mathcal{A}$ be an algorithm that w.h.p. in $\eta$ solves wakeup with $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$ latency, where $d^{\prime}>0$ is some constant. By Lemma 5.1, $\mathcal{A}$ has total contention:
(6) $\sum_{t=1}^{L}\texttt{Con}(t)\geq 2$
As we discussed, Lemma 5.2 implies that we need only consider that $\mathcal{A}$ has contention at most $2$ in every slot. Thus, we can use Lemma 5.3 (noting that $n\geq 2$), which implies that the expected cost of collisions in a single slot $t$ is:
@@
$\sum_{t=1}^{L}k\,\texttt{Con}^{2}(t)\,\mathcal{C}.$
Using Jensen’s inequality for convex functions, we have:
$\mathcal{C}k\sum_{t=1}^{L}\texttt{Con}^{2}(t)$ $\geq\mathcal{C}k\frac{\left(\sum_{t=1}^{L}\texttt{Con}(t)\right)^{2}}{L}$
$\geq\frac{\mathcal{C}\,k\,2^{2}}{L}$
$=\Omega(\mathcal{C}^{(1/2)-\alpha\epsilon}/\log^{\beta}n)$
where the second line follows from Equation 6, and the last line follows since $L\leq d^{\prime}\mathcal{C}^{(1/2)+\alpha\epsilon}\log^{\beta}n$. ∎
@@
6. Conclusion and Future Work
Our work gives a preliminary theoretical exploration addressing collision costs. For future work, we aim to prove more general lower bounds. We also wish to derive results for the dynamic setting, where packets arrive over time rather than all being initially present.
Acknowledgements. This research is supported by the National Science Foundation grants CNS-2210300 and CCF-2144410.